• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name jpchen

    Log entry time 16:14:42 on July15,2003

    Entry number 103873

    keyword=shift summary

    Day shift: shift workers: X. Zheng, E. Voutier, J. P. Chen.

    8:00 Beam tuning: clean up orbit and BCM calibration for beam loss monitoring
    10:00 Beam arrived. MCC tested to 20 uA, had raster working.
    10:30 Centering beam, fast feedback on, harp scan (2104).
    11:00 right Q2 tripped controlled access to reset.
    11:30 another harp scan (2105). x too large (~400 um), y divergence too much
    (40 um --> 150 um). OTR shows 160x150 um.
    12:30 MCC do multi-harp scan
    15:00 they completed multi-harp scan, did adjustments, we did a few harp to check beam spot size
    16:00 still adjusting beam spot size with harp scan help.